

********************************************************************
UARS WEEKLY STATUS REPORT - 17 DEC 1999
********************************************************************
***********
SOLSTICE
***********
Current tape recorder operations caused us about a 40% loss of our planned
experiments this week. Otherwise, SOLSTICE continues to operate normally,
collecting solar data on all available orbital tracking periods and stellar
calibration observations, whenever possible.
***********
SUSIM
***********
Nothing receievd.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM spectrometers continue to be active under PMON control*. While
under PMON control, the PEM spectrometers computed energy deposition from
electrons and protons measured in-situ by the particle spectrometers.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing receievd.
***********
WINDII
***********
Nothing receievd.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations. The
Latest Instrument Operations schedule covers the period
6 January to 15 February 2000.
The WINDII instrument team asked if the UARS spacecraft could
operate with a roll offset of 0.4 degrees while WINDII is
taking data in December. The MPG contacted other instrument
teams and the FOT. It was determined that WINDII's request
could be met. Roll offset operation was started on
15 December and will end on 23 December.
***********
CDHF
***********
The CDHF continues to notify all PI Software Representatives when Data Gaps
occur in the LEVEL0 data due to the UARS Spacecraft Anomaly.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
17 days CLA (7.0) PROD (L1-L3) (476R-452C-24F=0L)
21 days GPI (4.05) PROD (CLAES L3B) (397R-377C-3F=17L)
42 days WIN (5.11A) PROD (L1-L3) (520R-144C-6F=370L)
Completed the CLAES V7.0 mission reprocessing for a total
of 452 days. CLAES PI software representative has been notified.
UARS CDHF REHOST:
-Reprocessed HALOE V4.82 on the Rehost System. HALOE validated by Pat Purcell
in timely manner. Much appreciated.
-Reprocessed SUSIM V7.1 on the Rehost System. SUSIM has been notified output
files are ready for validation.
-Testing Sensor Extraction V5.0 software on the Rehost System.
-Testing Observatory Log V3.4 on the Rehost System.
-Testing SOLSTICE V4.62 on the Rehost System.
-Migration for all data levels of data via 8MM tape has been put on hold to
free resources for Production testing and migration and performance evaluation.
-Completed the configuration of the NAG (Numerical Algorithms Group) software.
-Installed IDL 5.0 on the rehost system in "demo" mode. For cost saving
purposes the license is to be transferred to the rehost system once it is
officially deemed operational.
-Continue to perform reboots of CDHFA/B as scheduled on Wednesdays to check
system settings, logicals, etc at startup as changes continue to be required
on these systems.
DAAC DATA:
-DAAC data through September 1999 is available to the User Community for
HALOE, PEM, NMC, and UKMO. October 1999 current day DAAC data will be loaded
in December 1999.
-SUSIM 3BS Version 20 from 1991-10-11 to 1998-08-08 is available at the DAAC.
***********
FOT
***********
REPORTING PERIOD
This report covers 4 December, 1999 (Orbit 44975, GMT Day 338) through 10
December 1999 (Orbit 45079, GMT Day 344).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 339/2339 45004 Off
The beta angle ranged between the angles of 21.4 degrees to 6.1 degrees for
this report period. The beta angle is now increasing toward a maximum angle of
+34.0 degrees on 23 December 1999 (DOY 357).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.0 to 1.3 degrees C. Additional battery performance data for each
battery on 03 December 1999 and 10 December 1999 are provided in Appendix D.
The clock error ranged from -10.9 to 10.9 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% OUT-OF-LIMITS = 04
% ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 02
% HALOE = 01
% HRDI = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
No Anomaly Reports initiated this report period.
Two (2) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
% Transponder A: 99/320 - 2287.496800
% Transponder B: 99/334 - 2287.499196
GROUND SYSTEM OPERATION
zero (0) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 99/343 show the observatory data loss to be 272 hours, 51
minutes, 08.537 seconds (an increase of 46 hours 19 minutes 59.584 seconds
since the last report period). This is a 0.3790 percent data loss which equals
a 99.6210 percent data capture for the mission.
The increase of 46 hours 19 minutes 59.584 seconds occurred over seven (7)
days:
- 06 hours 04 minutes 34.368 seconds on Day 337 due to a S/C TR problem.
- 05 hours 28 minutes 58.528 seconds on Day 338 due to a S/C TR problem.
- 06 hours 59 minutes 54.560 seconds on Day 339 due to a S/C TR problem.
- 07 hours 00 minutes 55.456 seconds on Day 340 due to a S/C TR problem.
- 07 hours 45 minutes 15.104 seconds on Day 341 due to a S/C TR problem.
- 06 hours 36 minutes 36.928 seconds on Day 342 due to a S/C TR problem.
- 06 hours 23 minutes 44.640 seconds on Day 343 due to a S/C TR problem.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
We also planned and implemented the HALOE / NBTR-B stored command test. This
test involved running NBTR-B out of the stored command load to catch all of
HALOE's sunrise and sunset events. This worked well and the real data
collection will begin on December 12-15.
The FOT gathered all the pertinent information about the tape recorder and sent
it up to Camden for review by Tim Ohrman and his group. Their company
manufactured these recorders and may have some insight into the problem.
The FOT prepared a data package for the folks that are recommending a new
trending system for UARS. A new trending system would reduce data processing
from hours to minutes or seconds.
WINDII sent in a request for a roll offset maneuver of +0.4 degrees for the
15th-23rd. FDF was contacted to see if they could get it into their planning
aid in time. They indicated that they could be ready, therefore preparations
are in progress for this event.
YEAR 2000 OPERABILITY SUPPORT
All areas identified to or by the Flight Operations Team for Y2K testing have
been completed. There are no other areas identified at this time that have not
been Y2K certified.
MEETINGS
Wednesday, December 8, 1999 (DOY 342) - UARS monthly meeting was held. The
NBTR status was discussed along with the resolution activities. Also discussed
were the critical science mission requirements. A splinter meeting followed to
discuss the future responsibility of the flight software.
FUTURE SIGNIFICANT EVENTS
No specific date was agreed upon for UARS Safehold configuration discussions.
Instead, the Supporting Engineers involved will be contacted every two weeks to
determine if either the Flight Operations Team or Supporting Engineers have any
new developments that should be discussed in a group.
UARS has tested and prepared for the Year 2000 event. Special considerations
have been made in preparation for a non nominal 1999 to 2000 transition.
The next UARS Yaw slew is scheduled for January 6, 2000. This maneuver will be
a Forward to Reverse Flight maneuver (Northern Hemisphere viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 10 December 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
44982 338/1237 99-409 OBC Clk Adj 0.00 to -20.6 msec/day
44996 339/1058 99-410 OBC Clk Adj -20.6 to 0.00 msec/day
45004 339/2339 99-411 HALOE Power OFF
45033 341/2203 99-412 HRDI Additional Microload
____________________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 10 December 1999
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPTSPBIMON Y-HI PTS PB Current Monitor 44974-5
HRPSEQBIMON Y-HI PSE QB Current Monitor 44994-5, 45041
HRPTSPBIMON R-LO HRDI Tele Barrel Curr. Mon. 45009
HRPTSPBIMON R-LO HRDI Tele Barrel Curr. Mon. 45024
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
No Alerts to report this report period.
____________________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 10 December 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
44982 338/12:50 (04 Dec 99) from 0.00 to -20.6 msec/day
44996 339/11:12 (05 Dec 99) from -20.6 to 0.00 msec/day
____________________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 10 December 1999
03 Dec 1999 - GMT Day 337 Beta = 21.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ---------- ------- -------- ----- ------- ------
1 -50.4/-72.8 -3.05/-4.62 +0.0/+0.0 16.2 0.000 0.0
2 +56.0/-5.6 +6.44/+5.15 +31.6/-15.2 24.8 1.037 20.0
3 +33.6/-33.6 +5.15/+3.87 +31.2/-16.0 24.5 1.035 19.7
10 Dec 1999 - GMT Day 344 Beta = 6.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ---------- ------- --------- ----- ------ ------
1 -44.8/-72.8 -3.68/-4.62 +0.0/+0.0 16.2 0.000 0.0
2 +16.8/-44.8 +5.15/+3.87 +30.4/-13.2 25.8 1.099 24.8
3 +39.2/-11.2 +3.87/+2.60 +29.6/-13.2 25.8 1.087 24.5
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- --- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 10 December 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
No new reports generated
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
No anomaly closures generated
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
No new reports generated
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-064 FOT 44135 10/8-9/99 N/A HRDI Cal Lamp Errors
99-071 FOT 44398 299/1212 21512 TPF Not Receiving 512 PB
SEVNTFW99/50